OpenStudy (anonymous):

@ajaylounsber Jennifer has been saving for college for 57 months. The first month, she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month?

OpenStudy (anonymous):

theres missing info

OpenStudy (anonymous):

$11 $11.50 $12 $12.50

OpenStudy (anonymous):

tell me more like how much needs to save up to ok?

OpenStudy (anonymous):

That's all they gave me.

OpenStudy (anonymous):

jeremyggg take it awy

OpenStudy (anonymous):

Did you type it exactly word for word? It does seem to be worded strange.

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

yes they are word for word

OpenStudy (anonymous):

Im thinking it's a set amount more each month. 11 22 33 44 55 type thing ajay

OpenStudy (anonymous):

19,768

OpenStudy (anonymous):

so if it is a set so how do i know how much... to be honest the question is so weird to me

OpenStudy (anonymous):

the answer i typed in is right

OpenStudy (anonymous):

Which class is this for? There could be a few ways to solve this

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

im taking it online

OpenStudy (anonymous):

simple subtrsction problem

OpenStudy (anonymous):

lily have you learned about factorials and summations in this class? does what i showed you look familar?

OpenStudy (anonymous):

$11 $11.50 $12 $12.50 it has to be one of them

OpenStudy (anonymous):

Yes it does.

OpenStudy (anonymous):

add to 11 57 times

OpenStudy (anonymous):

one of those answers

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wording is confusing because if she was able to save more money then the month before, how can what she saved be a constant rate

OpenStudy (anonymous):

She saves an exact amount more than the previous. like a series or summation. 11 + (11+x) + (11+ 2x)+(11+3x) ....(11+57x) = 19779 do you know how to set this up in a summation lilystacy? maybe you can give it a try?

OpenStudy (anonymous):

No i do not know how.....

OpenStudy (anonymous):

Ok well, when you have a summation (this symbol)\[\sum_{k}^{n}\] you add these values over and over where k is your starting point and n is the end point. we know the first month needs to be 11 + 0 and each additional month is adding 1 to the multiplier of x. sounds confusing i know. but this is what we are doing. month 1: 11 + 0x month 2: 11+ 1x (because x is the amount she saves more than last month) month 3: (11+ 1x) + x (11 +1x is the amount she saved last month) month 4 (11+2x)+x and so on. we have our pattern! so we want to start at 0 so that we can get 11. if we start at zero we need to go to 56 to make a total of 57 months. \[\sum_{k=0}^{56} 11 +kx = 19779\] does your calculator let you type this to solve for x?

OpenStudy (anonymous):

i have my calculator right now but i do not know how to plug them in

OpenStudy (anonymous):

Hmm, i am not sure of your calculator. I know a website I use called wolframalpha.com where i sometimes checked my answers when i forgot my calculator. The answer I got when plugging this in is 12.

OpenStudy (anonymous):

yayyyy Jeremyggg it's right you did it

OpenStudy (anonymous):

and this is what i handtyped in the website. summation k =0 to 56 (11 + xk) = 19779 those exact words!

OpenStudy (anonymous):

There is a hand way of doing this. I'm trying to remember the steps so maybe you can do these without a calculator! do you think you need these steps?

OpenStudy (anonymous):

Can you help me again please given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain. n an -4 20 -100 1 2 3

OpenStudy (anonymous):

an = −5(−4)n − 1 where n ≥ 1 an = −4(−5)n − 1 where n ≥ 1 an = −4(5)n − 1 where n ≥ −4 an = 5(−4)n − 1 where n ≥ −4

OpenStudy (anonymous):

ooh i used to like these i think. ok let's see...

OpenStudy (anonymous):

There is a formula we need. something along the lines of \[a _{n} = a _{1}*r ^{n-1}\] does this one sound familiar or right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where r is the ratio between each number

OpenStudy (anonymous):

ok we have a1 already right? now let's figure out r! (let me know if/when you get lost)

OpenStudy (anonymous):

I'm here

OpenStudy (anonymous):

ok, just to make sure you are keeping up, let me know what a1 and r equal if you know em! they are simple, but this stuff can get confusing really fast.

OpenStudy (anonymous):

This is the first time i'm seeing those thing and they gave me a lesson i don't even know what they are saying,,,,

OpenStudy (anonymous):

Jeremyggg are you there

OpenStudy (anonymous):

yes sorry!

OpenStudy (anonymous):

alright well, let's teach how this works. n an 1 -4 2 20 3 -100 this is saying that we have this series (i like to think of em as patterns or puzzles) where a1 = -4 a2= 20 and a3=-100 r is going to be a ratio between them, and being a geometric sequence, this ratio will be constant. what do you do to a1 to make a2=20 and what do you do to a2 to make a3= -100? this is your r!

OpenStudy (anonymous):

with this info we have most of our answer already! an = a1* r^n-1 or... an= -4(r)^n-1

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

Yes, but what is r?

OpenStudy (anonymous):

remember r as the word ratio!

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

i guess it might help that i explain what we are doing. we are coming up with an equation where if i wanted to know what a10 was, i could just plug 10 into the equation that we created. like for example, what we solved earlier! i wanna know the 57th month, i just plug 57 into the equation and im done!

OpenStudy (anonymous):

no no. let us review the table again. a1= -4 a2 = 20 a3= -100 r = the ratio between each of these numbers. there is something that happens to a1 that makes it equal a2. the exact same thing happens to a2 to make it a3. and if there was an a4, the exact same thing would happen to 3 to make a4.

OpenStudy (anonymous):

So 4 is wrong then we only have 1

OpenStudy (anonymous):

the ratio is -5! -4 * -5 = 20 20*-5 = 100! so r = -5

OpenStudy (anonymous):

to go from a1 to a2 we must multiply by -5. to go from a2 to a3 we must multiply by -5. this is what i am meaning by ratio

OpenStudy (anonymous):

It can't be -5 remember we have an = −5(−4)n − 1 where n ≥ 1 an = −4(−5)n − 1 where n ≥ 1 an = −4(5)n − 1 where n ≥ −4 an = 5(−4)n − 1 where n ≥ −4 that's are the options they gave me

OpenStudy (anonymous):

yes. and remember the equation i gave you? an = a1*r^n-1

OpenStudy (anonymous):

a1 = -4 based on the chart. r = -5 based on what we solved for the ratio! an = a1*r^n-1 orrrrr.... an = -4(-5)^n-1

OpenStudy (anonymous):

B IS THE ANSWER

OpenStudy (anonymous):

yes! The bounds can be a bit tricky to explain until you grasp what is going on with this one a bit better

OpenStudy (anonymous):

erm... restrictions on domain, not bounds

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

No problem! If you need any more help just send me a message! And if you need a better explaining than this, because this was one of my worst attempts at teaching, also let me know!

OpenStudy (anonymous):

Alright right now i'm looking at what i need help with if i find some i let you know.

OpenStudy (anonymous):

just now finish it wow

OpenStudy (anonymous):

ooh we did 2 of them ajay!

OpenStudy (anonymous):

oh ok but dang

OpenStudy (anonymous):

next time don't use one with some one tagged in the question it get's anoying

OpenStudy (anonymous):

you left us

OpenStudy (anonymous):

I said for him to take it away

OpenStudy (anonymous):

Oh Ok

OpenStudy (anonymous):

A study of college freshmen found that students who watched TV for more than one hour each night gained 15 pounds in the first semester, compared to those who watched less than one hour. They concluded that there is an association between watching TV and weight gain in college freshmen. Which study design did they most likely use?

OpenStudy (anonymous):

Answers They most likely used an experiment with students randomly assigned to watch more than one hour or less than one hour of TV each night. They most likely used an experiment with students randomly assigned to groups based on how much weight they gained. They most likely used an observational study asking students to keep track of their TV habits and weight. They most likely used an observational study asking students to keep track of their workout regimen and weight.

OpenStudy (anonymous):

Jeremyggg are you there i have five more 3 of them i have the answers and i don't know 2 of them could you help me please.

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

use diffrent onr please